# ESTIMATION OF FIRST RETURNING SPEED IN NULL RECURRENT CONTINUOUS-TIME MARKOV CHAINS

### v1

@@ 5.1 Linear growth model
According to the Raabe’s Ratio Test, the series $\sum_{i=0}^{\infty}a_{i}^{-1}$ diverges when $s-r\leq m$, and converges when sr>ms-r>m. That is to say, $\{X_{t},t\geq 0\}$ is recurrent if and only if $k>m$ or $s-r\leq k=m$.
@@
Let $w(t)=\log^{\gamma}t$ and $\gamma>0$. By Corollary 5.2, $E(\log^{\gamma}\tau)<\infty$ if and only $\int^{\infty}|w^{\prime\prime}(t)|F^{-1}(t)dt<\infty,$ where $F(n)\sim n\log n.$ Based on the asymptotic estimation of the Lambert $W$-function, we can obtain $F^{-1}(t)\sim\dfrac{t}{\log t}.$ It follows that
$\int^{\infty}|w^{\prime\prime}(t)|F^{-1}(t)dt\sim\int^{\infty}\dfrac{1}{t\log^{2-\gamma}t}dt.$
Therefore, $E(\log^{\gamma}\tau)<\infty$ if and only if $0<\gamma<1$.
@@
Take $w(t)=t^{a}$ and $0<a<1$. Similarly, by Corollary 5.2, we have
@@
Therefore, $E(\tau^{a})<\infty$ if and only if $0<a<1-\dfrac{s-r}{m}$. ∎
@@ 5.2 Bilateral birth-death process
$\left\{\begin{aligned} q_{ij}&=0(|i-j|>1)\\ q_{ii-1}&=\alpha_{i}>0\\ q_{ii+1}&=\beta_{i}>0\\ q_{i}:=&-q_{ii}=\alpha_{i}+\beta_{i}<\infty(\forall i\in\mathbf{Z}).\end{aligned}\right$.
@@
$\left\{\begin{aligned} m_{n}&=\dfrac{\alpha_{-1}\alpha_{-2}\cdots\alpha_{n+1}}{\beta_{0}\beta_{-1}\cdots\beta_{n+1}\beta_{n}}(n<-1),\\ m_{-1}&=\dfrac{1}{\beta_{0}\beta_{-1}},\,m_{0}=\dfrac{1}{\alpha_{0}\beta_{0}},\,m_{1}=\dfrac{1}{\alpha_{0}\alpha_{1}},\\ m_{n}&=\dfrac{\beta_{1}\beta_{2}\cdots\beta_{n-1}}{\alpha_{0}\alpha_{1}\alpha_{2}\cdots\alpha_{n-1}\alpha_{n}}(n\geq 1).\end{aligned}\right$.
@@
It is easy to obtain that the invariant distribution of the process is constant, which follows $a_{n}=\beta_{n},b_{n}=\lambda$. Then we have $A^{+}_{2N}\sim A^{-}_{2N}\sim\dfrac{rs}{N(r+s)},B_{N}^{+}(\lambda)\sim B_{N}^{-}(\lambda)\sim\lambda N.$ Take $N=\sqrt{\dfrac{rs}{\lambda(r+s)}}$, by Theorem 4.9, we have $C_{0\leftrightarrow\Delta}(\lambda)\sim\sqrt{\lambda}.$ According to Theorem 4.4, assume that $w(t)=t^{a}$ where $0<a<1$, we have
@@
Hence $E(\tau^{a})<\infty$ if and only if $0<a<\dfrac{1}{2}$. ∎
@@
$\pi(i)=\dfrac{3}{2^{|i|+3}}(i\neq 0),\,\pi(0)=\dfrac{1}{4},$
@@
    If $\gamma=2$, $A_{N}^{+}\sim A_{N}^{-}\sim 2^{-N}$ and $B_{N}^{+}(\lambda)\sim B_{N}^{-}(\lambda)\sim\lambda N$. Take $N=\exp[W(1/\lambda\log 2)]\sim(-\lambda\log\lambda)^{-1}$, where $W(x)$ is Lambert W-function. Then we have
    $C_{0\leftrightarrow\Delta}(1/t)\sim\dfrac{1}{\log t}.$
    Assume that $w(t)=\log^{\gamma}t$ and $\gamma>0$, according to Theorem 4.4, we have
    $tw^{\prime\prime}(t)C_{0\leftrightarrow\Delta}(1/t)\sim(t\log^{2-\gamma}t)^{-1}.$
    Hence $E(\log^{\gamma}\tau)<\infty$ if and only if $0<\gamma<1$.
  2. (2)
@@
    Hence $E(\tau^{a})<\infty$ if and only if $0<a<\log_{\gamma}2$.

### v2

@@ 5.1 Linear growth model
According to the Raabe’s Ratio Testhuynh2021secondraabestestseries, the series $\sum_{i=0}^{\infty}a_{i}^{-1}$ diverges when $s-r\leq m$, and converges when sr>ms-r>m. That is to say, $\{X_{t},t\geq 0\}$ is recurrent if and only if $k>m$ or $s-r\leq k=m$.
@@
Let $w(t)=\log^{\gamma}t(\gamma>0)$. Based on the asymptotic estimation of the Lambert $W$-function, we can obtain $F^{-1}(t)\sim\dfrac{t}{\log t}.$ By Theorem 5.1, we have
$\int^{\infty}|w^{\prime\prime}(t)|M(F^{-1}(t))dt\sim\int^{\infty}\dfrac{1}{t\log^{2-\gamma}t}dt.$
Therefore, for $\gamma>0$, $E(\log^{\gamma}\tau)<\infty$ if and only if $\gamma<1$.
@@
Take $w(t)=t^{a}(0<a<1)$. Similarly, by Theorem 5.1, we have
@@
Therefore, for $a>0$, $E(\tau^{a})<\infty$ if and only if $a<1-\dfrac{s-r}{m}$. ∎
@@ 5.2 Bilateral birth-death process
$\left\{\begin{aligned} q_{ij}&=0,\qquad|i-j|>1,\\ q_{i,i-1}&=\alpha_{i}>0,\\ q_{i,i+1}&=\beta_{i}>0,\\ q_{i}:=&-q_{ii}=\alpha_{i}+\beta_{i}<\infty.\end{aligned}\right$.
@@
$\left\{\begin{aligned} m_{n}&=\dfrac{\alpha_{-1}\alpha_{-2}\cdots\alpha_{n+1}}{\beta_{0}\beta_{-1}\cdots\beta_{n+1}\beta_{n}},\,\,n<-1,\\ m_{-1}&=\dfrac{1}{\beta_{0}\beta_{-1}},\,m_{0}=\dfrac{1}{\alpha_{0}\beta_{0}},\,m_{1}=\dfrac{1}{\alpha_{0}\alpha_{1}},\\ m_{n}&=\dfrac{\beta_{1}\beta_{2}\cdots\beta_{n-1}}{\alpha_{0}\alpha_{1}\alpha_{2}\cdots\alpha_{n-1}\alpha_{n}},\,\,n>1.\end{aligned}\right$.
@@
It is easy to obtain that the invariant measure $m_{i}=\dfrac{1}{rs}$ for $i\in\mathbf{Z}$. It follows that $A^{+}_{N}\sim A^{-}_{N}\sim\dfrac{1}{N},B_{N}^{+}(\lambda)\sim B_{N}^{-}(\lambda)\sim\lambda N.$ Take $N=\sqrt{\dfrac{1}{\lambda}}$, by Theorem 4.9, we have $C_{0\leftrightarrow\Delta}(\lambda)\sim\sqrt{\lambda}.$ According to Theorem 4.4, assume that $w(t)=t^{a}$ where $0<a<1$, we have
@@
Hence for $0<a<1$, we have $E(\tau^{a})<\infty$ if and only if $0<a<\dfrac{1}{2}$. ∎
@@
$\pi(i)=\dfrac{3}{2^{|i|+3}}\mbox{\,\,for\,\,}i\neq 0,\quad\,\pi(0)=\dfrac{1}{4},$
@@
    Hence, for $0<a<1$, $E(\tau^{a})<\infty$ if and only if $0<a<\log_{\gamma}2$.
  2. (2)
    If $\gamma=2$, $A_{N}^{+}\sim A_{N}^{-}\sim 2^{-N}$ and $B_{N}^{+}(\lambda)\sim B_{N}^{-}(\lambda)\sim\lambda N$. Take $N=\left[\dfrac{W(\frac{\log 2}{\lambda})}{\log 2}\right]\sim\log\dfrac{1}{\lambda}$, where $W(x)$ is Lambert W-function. Then we have
    $C_{0\leftrightarrow\Delta}(1/t)\sim\dfrac{\log t}{t}.$
    Let $w(t)=t^{a}(0<a<1)$, according to Theorem 4.4, we have
    $t|w^{\prime\prime}(t)|C_{0\leftrightarrow\Delta}(1/t)\sim\dfrac{\log t}{t^{2-a}}.$
    Hence $E(\tau^{a})<\infty$ always holds for $0<a<1$.
    If we take $w(t)=\dfrac{t}{\log^{b}t}(b>0)$, then we have
    $t|w^{\prime\prime}(t)|C_{0\leftrightarrow\Delta}(1/t)\sim\dfrac{1}{t\log^{b}t}.$
    Hence, for $b>0$, $E(w(\tau))<\infty$ always holds for $b>1$.